OpenStudy (anonymous):

Which of the following shows the correct value of the ? 14 < < 15 17 < < 18 18 < < 19 21 < < 22

OpenStudy (adamaero):

what?

OpenStudy (anonymous):

The space between the numbers are 175 with a check mark thing

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

i am going to guess that it is something like \[14<\sqrt{175}<15\] and so on am i right?

OpenStudy (anonymous):

Yes you're right @misty1212

OpenStudy (misty1212):

ok then we can do it

OpenStudy (anonymous):

I honestly think it is a..

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

oh wait that is not right

OpenStudy (misty1212):

none of the answers are right

OpenStudy (misty1212):

\[14^2=196\] which is bigger than \(175\)

OpenStudy (anonymous):

hmm, well dang...

OpenStudy (misty1212):

should be\[13<\sqrt{175}<14\]

OpenStudy (anonymous):

Alright, thanks(:

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]